# Understanding and Estimating the Execution Time of Quantum Programs

### v1

@@ 6.1.3 Evaluation of the accuracy of execution times estimated by the IBM Quantum Platform
The estimation accuracy of the IBM Quantum Platform for execution times is notably inadequate. By observing the MSE, R-squared, and NMSE, as shown in Table 5, we realize that the IBM Quantum Platform’s predictive performance for execution times is very poor. For example, the $R^{2}$ value is negative, indicating that the prediction results are less accurate than simply predicting the average execution time. This necessitates the proposal of a new approach to better estimate the execution times of quantum circuits on quantum computers.
Highlights of RQ1 On both simulators and real quantum computers, the execution times of quantum circuits vary significantly across different circuits, yet the execution time estimation provided by IBM Quantum is far from being accurate, highlighting the need for an approach to accurately predict the execution times of quantum circuits.
6.2 RQ2: How well can we estimate the execution times of quantum circuits on simulators, and what are the important predictors?
@@
6.2.1 The overall performance of the model in predicting the execution times of quantum circuits on simulators
Based on the evaluation method and metrics described in 5.4, in a 10-fold cross-validation, we evaluate the performance of our model for predicting the execution times of the quantum circuits on simulators. Table 6 shows the evaluation results.
Table 6: Evaluation of our model’s prediction results on simulators
Data Points: 3020 Simulators: FakeWashington, FakeSherbrooke
@@
6.2.2 Assessment of the importance of different components of the graph features
@@
Table 7: Evaluation of prediction results on simulators with missing components of the graph features
Data Points: 3020 Simulators: FakeWashington, FakeSherbrooke
@@
6.2.3 Assessment of the importance of different components of the global features
@@
SHAP values quantify the contribution of each feature to a prediction based on Shapley values from game theory, ensuring a fair distribution of feature importance. The core idea of SHAP values is to measure the impact of a feature on the model’s prediction by considering the difference in the prediction when the feature is present versus when it is absent. By calculating the contribution of each feature value and averaging these contributions across all possible combinations of features, SHAP values provide a fair measure of each feature’s importance [12]. A positive SHAP value indicates that the feature increases the predicted value, whereas a negative SHAP value suggests that the feature decreases the predicted value. The larger the absolute value of a SHAP value, the greater its impact on the model’s predicted value, and correspondingly, the higher the contribution of the corresponding feature to the model. Therefore, on the test set, for each feature in the global features, we calculate the average of the absolute values of the SHAP values predicted for each data point. These averages are then sorted from largest to smallest, with larger values indicating more important features. The results are visualized in Fig. 9.
@@
Figure 9: The mean of the absolute values of the SHAP values for different features of the global features on simulators
The number of gates, particularly two-qubit gates such as the Controlled-NOT and Controlled-Phase gates, plays the most important roles among the global features in predicting the execution times of quantum circuits. As shown in Fig. 9, due to the large number of features in the global features (41), we only display the top 10 most important features. From Fig. 9, it can be observed that, apart from the depth and number of qubits features, the other important features are various quantum gates. The two most significant gates are the Controlled-NOT and Controlled-Phase gates, both of which are two-qubit gates. These gates are very common in quantum circuits, and it can be inferred that their prevalence also reflects the complexity of a quantum circuit, thereby playing a significant role in predicting the execution time of the circuit. In particular, two-qubit gates are known to have longer gate time (the time taken for a gate operation) than one-qubit gates [28].
@@
Highlights of RQ2 Our graph transformer-based model, based on the structure (graph) and global features of quantum circuits, can accurately predict the execution times of quantum circuits on simulators, achieving an $R^{2}$ score of 0.943 (the $R^{2}$ further increases to 0.961 when the Node Type feature is excluded). Overall, the graph features play a more important role than the global features; among the individual features, the T1, T2 features within the graph features, and the number of different types of two-qubit gates (e.g., controlled-not gates) within the global features, are among the most important features.
@@
6.3 RQ3: How well can we estimate the execution times of quantum circuits on real quantum computers, and what are the important predictors?
@@
Similar to RQ2 (6.2), we build and evaluate our model for predicting the execution times of the quantum circuits on real quantum computers, based on the approach described in 5. Same as in RQ1, we consider the 340 selected circuits that are selected to execute on real computers: 192 circuits on ibm_osaka and 148 on ibm_kyoto. The details for the circuit selection through active learning are described in 5.3.
@@
6.3.1 The overall performance of the model in predicting the execution times of quantum circuits on quantum computers
@@
Table 8: Evaluation of prediction results on quantum computers
@@
6.3.2 Assessment of the importance of different components of the graph features
@@
All features within the graph features positively impact the model’s performance; however, none of the graph features are critical, as removing any one of these features does not significantly impact the model’s performance. In particular, unlike the results of RQ2, the T1, T2 features do not play a significantly important role: in fact, the absence of the T1, T2 features results in the smallest increases in Average MSE and Average NMSE, and the least decrease in Average R-squared, which contrasts sharply with their highest importance for predicting execution times on simulators in RQ2. As explained in RQ2, a longer T1 or T2 time may require a simulator to take a longer time to simulate the circuit behavior; however, when a circuit is executed in a real quantum computer, intuitively, the execution time does not depend on the T1, T2 times: T1, T2 times control the maximum time a quantum circuit can execute reliably, not the actual execution time. Our observations highlight the intrinsic differences between modeling quantum circuits’ execution times on real quantum computers and simulators.
Table 9: Evaluation of prediction results on quantum computers with missing components of the graph features
@@
6.3.3 Assessment of the importance of different components of the global features
@@
Figure 10: The mean of the absolute values of the SHAP values for different features of the global features on quantum computers
The numbers of different types of gates, particularly two-qubit gates such as Controlled-Phase and Controlled-NOT gates, remain the most crucial features among the global features in predicting the execution times of quantum circuits on quantum computers. We find that, compared to those identified in the model on the simulators, the top 10 most important features among the global features on quantum computers lack the number of qubits feature, but include the rotation-y gate feature. The remaining features remain the same. This indicates that, whether on simulators or on quantum computers, the quantity of common gates within quantum circuits plays a crucial role in accurately predicting the execution times of these circuits. In particular, despite the swapping of their rankings, the numbers of two-qubit gates (e.g., control-phase gate, controlled-not gate) still play the most important roles, due to the fact that two-qubit gates have significantly longer gate times than single-qubit gates [28].
@@
7 Threats to Validity
7.1 Internal Validity
When executing quantum programs on quantum computers, the scarcity of quantum computer resources and the lengthy total time required for executing each circuit (including queuing time which usually takes hours for a circuit), prevent us from executing all the quantum programs on quantum computers. Instead, we selected a subset of the quantum programs with 95% confidence for execution and data collection. We used active learning to maximize the diversity of teh sampled programs. However, this may result in a deviation between the observed results and the actual results: for example, the performance of our model may increase if more quantum circuits are considered.
7.2 External Validity
@@ 7.3 Construct validity
8 Conclusion
Quantum computing resources are rather limited, thus making judicious planning and utilization of quantum computers is important for both quantum computing service providers and users. In this paper, we first study the characteristics of quantum programs’ runtime on simulators and real quantum computers. Then, we introduce an innovative graph transformer-based model, utilizing the graph information and global information of quantum circuits to estimate their execution times. Our experiments indicate that the execution times of quantum circuits vary significantly across different circuits, and that our approach can accurately predict their execution times on both simulators and quantum computers. We observe that some characteristics of the quantum circuits, such as the numbers of different types of quantum gates, especially two-qubit gates, play important roles in predicting their execution times, yet we also highlight the intrinsic differences in modeling the execution times of quantum circuits executed on simulators and real computers.
@@ 8 Conclusion
Our study also sheds light on further explorations. First, future research may leverage our approach to automatically optimize quantum circuits in terms of their execution time consumption.
In addtion, future research may extend our study beyond the IBM Quantum Platform to include other emerging quantum computing technologies, such as Google’s Sycamore or Rigetti, which could provide broader insights for understanding quantum execution times and our model’s adaptability and efficiency across different quantum systems.

### v2

@@ 6.1.3. Evaluation of the accuracy of execution times estimated by the IBM Quantum Platform
The estimation accuracy of the IBM Quantum Platform for execution times is notably inadequate. By observing the MSE, R-squared, and NMSE, as shown in Table 9, we realize that the IBM Quantum Platform’s predictive performance for execution times is very poor, since it reports a worst case circuit execution time rather than an expected value. For example, the $R^{2}$ value is negative, indicating that the prediction results are less accurate than simply predicting the average execution time. This necessitates the proposal of a new approach to better estimate the execution times of quantum circuits on quantum computers.
Highlights of RQ1 On both simulators and real quantum computers, execution times vary substantially across circuits. Simple, target-independent features have nontrivial but limited associations with execution time, often monotonic rather than linear. These observations underscore the need for learned models that combine multiple features, while the backend-to-backend differences in the associations motivate the consideration of backend-dependent features and models. Accordingly, RQ2 and RQ3 aim to develop accurate, backend-aware, model-based predictions.
6.2. RQ2: How well can we estimate the execution times of quantum circuits on simulators, and what are the important predictors?
@@
6.2.1. The overall performance of the model in predicting the execution times of quantum circuits on simulators
Based on the evaluation methods and metrics described in Section 5.4, we evaluate the performance of our model in predicting the execution times of quantum circuits on simulators using a test set comprising 10% of the entire dataset. The test set, with 302 data instances, is relatively large to provide a robustness evaluation of the model performance. Table 10 shows the evaluation results.
Table 10. Evaluation of our model’s prediction results on simulators
Test Data Points: 302 Simulators: FakeWashington, FakeSherbrooke
@@
6.2.2. Assessment of the importance of different components of the graph features
@@
Table 11. Evaluation of prediction results on simulators with missing components of the graph features
Test Data Points: 302 Simulators: FakeWashington, FakeSherbrooke
@@
The substantial impact of the T1 and T2 features on execution time prediction can be theoretically explained by the underlying simulation mechanism in Qiskit. When using a simulator such as FakeWashington or FakeSherbrooke, Qiskit automatically constructs a noise model based on the backend’s calibration snapshot [^11] https://qiskit.github.io/qiskit-aer/tutorials. This snapshot includes qubit-level properties such as T1 and T2, gate durations, and error rates. During simulation, T1 and T2 are used to model decoherence processes (e.g., amplitude and phase damping), which are inserted between operations or during idle time. Longer coherence times thus require the simulator to maintain precise density matrix evolution for a longer period, increasing the overall runtime. This explains the marked performance degradation observed when T1 and T2 features are excluded from the model.
6.2.3. Assessment of the importance of different components of the global features
@@
SHAP values quantify the contribution of each feature to a prediction based on Shapley values from game theory, ensuring a fair distribution of feature importance. The core idea of SHAP values is to measure the impact of a feature on the model’s prediction by considering the difference in the prediction when the feature is present versus when it is absent. By calculating the contribution of each feature value and averaging these contributions across all possible combinations of features, SHAP values provide a fair measure of each feature’s importance (lundberg2017unified). A positive SHAP value indicates that the feature increases the predicted value, whereas a negative SHAP value suggests that the feature decreases the predicted value. The larger the absolute value of a SHAP value, the greater its impact on the model’s predicted value, and correspondingly, the higher the contribution of the corresponding feature to the model. Therefore, on the test set, for each feature in the global features, we calculate the average of the absolute values of the SHAP values predicted for each data point. These averages are then sorted from largest to smallest, with larger values indicating more important features. The results are visualized in Fig. 9.
@@
Figure 9. The mean of the absolute values of the SHAP values for different features of the global features on simulators
The number of gates, particularly two-qubit gates such as the Controlled-Not (CX) and Controlled-Phase (CP) gates, plays the most important roles among the global features in predicting the execution times of quantum circuits. As shown in Fig. 9, due to the large number of features in the global features (41), we only display the top 10 most important features. From Fig. 9, it can be observed that, apart from the depth and number of qubits features, the other important features are various quantum gates. The two most significant gates are the CX and CP gates, both of which are two-qubit gates. This observation aligns with physical and compilation characteristics of quantum circuits. These gates are very common in quantum circuits, and it can be inferred that their prevalence also reflects the complexity of a quantum circuit, thereby playing a significant role in predicting the execution time of the circuit. Two-qubit gates are significantly more expensive to simulate than single-qubit gates because they can create entanglement between qubits, thereby increasing the complexity of the quantum state the simulator must track. Moreover, these two-qubit gates often undergo nontrivial decomposition during compilation. For example, a CP gate is typically decomposed into a combination of basis gates including multiple CX gates and single-qubit rotations. On FakeWashington, the CX gate is the native two-qubit gate; however, on FakeSherbrooke, the native entangling gate is ECR gate, meaning every CX gate in the circuit is further decomposed into multiple ECR gates and surrounding rotations. As a result, circuits containing high-level two-qubit gates—especially on non-CX-native devices—undergo deeper compilation and produce a larger number of physical-level operations for simulation. This explanation is quantitatively supported by Table 4, which lists average gate durations on both simulators. The native two-qubit gates, CX (550.41 ns) and ECR (533.83 ns), are nearly an order of magnitude slower than the single-qubit gates SX and X (35.56–56.89 ns). RZ gates incur almost no time cost. This stark difference explains why two-qubit gates dominate simulation time prediction: not only are they more expensive to simulate per gate, but they also propagate noise, entanglement, and depth, further burdening the simulator—especially when their number is inflated due to compilation and qubit topology constraints.
@@
Highlights of RQ2 Our graph transformer-based model, based on the structure (graph) and global features of quantum circuits, can accurately predict the execution times of quantum circuits on simulators, achieving an $R^{2}$ score of 0.943 (the $R^{2}$ further increases to 0.961 when the Node Type feature is excluded). Overall, the graph features play a more important role than the global features; among the individual features, the T1, T2 features within the graph features, and the number of different types of two-qubit gates (e.g., CX gates) within the global features, are among the most important features.
@@
6.3. RQ3: How well can we estimate the execution times of quantum circuits on real quantum computers, and what are the important predictors?
@@
Similar to RQ2 (6.2), we build and evaluate our model for predicting the execution times of the quantum circuits on real quantum computers, based on the approach described in 5. Same as in RQ1, we consider the 340 selected circuits that are selected to execute on real computers: 192 circuits on ibm_osaka and 148 on ibm_kyoto. The details for the circuit selection through active learning are described in 5.3.2.
@@
6.3.1. The overall performance of the model in predicting the execution times of quantum circuits on quantum computers
@@
Table 12. Evaluation of prediction results on quantum computers
@@
6.3.2. Assessment of the importance of different components of the graph features
@@
All features within the graph features positively impact the model’s performance; however, none of the graph features are critical, as removing any one of these features does not significantly impact the model’s performance. In particular, unlike the results of RQ2, the T1, T2 features do not play a significantly important role: in fact, the absence of the T1, T2 features results in the smallest increases in Average MSE and Average NMSE, and the least decrease in Average R-squared, which contrasts sharply with their highest importance for predicting execution times on simulators in RQ2. As explained in RQ2, a longer T1 or T2 time may require a simulator to take a longer time to simulate the circuit behavior; however, when a circuit is executed in a real quantum computer, intuitively, the execution time does not depend on the T1, T2 times: Unlike noise-aware simulators that explicitly simulate decoherence based on T1 and T2 values, real quantum computers only rely on pulse calibration and gate scheduling, where T1 and T2 represent the upper bounds on coherent operation windows, rather than determining the actual gate time. Therefore, T1 and T2 do not directly affect the execution time of a circuit, which instead depends on gate durations, pulse alignment, and scheduling constraints on hardware. This explains why removing the T1, T2 features does not significantly degrade performance in this setting. Our observations highlight the intrinsic differences between modeling quantum circuits’ execution times on real quantum computers and simulators.
Table 13. Evaluation of prediction results on quantum computers with missing components of the graph features
@@
6.3.3. Assessment of the importance of different components of the global features
@@
Figure 10. The mean of the absolute values of the SHAP values for different features of the global features on quantum computers
The numbers of different types of gates, particularly two-qubit gates such as CP and CX gates, remain the most crucial features among the global features in predicting the execution times of quantum circuits on quantum computers. We find that, compared to those identified in the model on the simulators, the top 10 most important features among the global features on quantum computers lack the number of qubits feature, but include the rotation-y gate feature. The remaining features remain the same. This indicates that, whether on simulators or on quantum computers, the quantity of common gates within quantum circuits plays a crucial role in accurately predicting the execution times of these circuits. In particular, despite the swapping of their rankings, the numbers of two-qubit gates (e.g., CP gate, CX gate) still play the most important roles.This is consistent with the compilation behavior on real quanutm devices. For example, both ibm_osaka and ibm_kyoto support only a limited native gate set, specifically [ecr, id, rz, sx, x], and lack native CX or CP support. As a result, gates like CX and CP must be decomposed into one or more ECR gates combined with single-qubit rotations. This decomposition increases the number of scheduled instructions and total circuit duration, leading to longer execution times—particularly for circuits with many high-level two-qubit gates.
Many previous studies have emphasized the importance of T gates in quantum cost analysis, such as (rahman2022grover). However, it is noteworthy that in this paper, the importance of the T gate is relatively low in both execution time estimation on simulators and on real quantum computers. Specifically, the T gate does not appear among the top-10 most important global features as determined by our feature importance analysis. Upon further inspection of the dataset used in this work, which consists of 1,510 quantum circuits, we find that only 104 out of 1,510 circuits contain any T gates, and the number of T gates in those circuits is generally small. This limited presence likely explains why the T gate does not have a significant impact on execution time prediction in our model. As a result, although the T gate may play a crucial role in other contexts, it does not contribute substantially to performance estimation in our setting.
@@
7. Discussion
In RQ3 (6.3), we build and evaluate our model for predicting the execution times of the quantum circuits on real quantum computers. Although the selected dataset contains quantum circuits with qubit counts ranging from 2 to 127, we note that circuits generated from the same algorithm family can still exhibit similar structural patterns, even if their qubit counts differ. Therefore, using a standard cross-fold validation on the dataset where many circuits originate from the same algorithm family may result in significant overlap between the training and test sets, potentially hindering the evaluation of the model’s generalization capability.
To address this, we propose a new cross-validation strategy that partitions the dataset by algorithm families. Specifically, in each fold, we select one algorithm family as the test set and use circuits from the remaining algorithm families as the training set. We observe that the selected 340 circuits span 10 distinct algorithm families; however, one algorithm family contains only two circuits. To ensure sufficient representation, we merge this algorithm family with the algorithm family containing the second fewest (26) circuits, resulting in a total of 9 folds for cross-validation.
Consequently, the number of circuits in each fold is [40, 46, 30, 35, 38, 28, 47, 40, 36]. The prediction performance obtained with this algorithm-family-based cross-validation scheme is summarised in Table 14.
Table 14. Prediction performance under the proposed algorithm-family-based cross-validation (9 folds)
Data Points: 340 Computers: ibm_osaka, ibm_kyoto
Model / Setting Average MSE Average R-squared Average NMSE
Full Features 0.485 0.878 0.122
Compared with the results in Section 6.3.2, the algorithm-family-partitioned cross-validation reveals notable differences: the average $R^{2}$ score falls from 0.905 to 0.878, the average MSE increases from 0.305 to 0.485, while the average NMSE rises from 0.095 to 0.122.
This indicates that when the training and test sets contain circuits originating from the same algorithm family, those circuits, despite large variations in qubit count, still share strong structural similarities that artificially improve prediction accuracy. Hence a random split yields an optimistic assessment of performance. After restricting each algorithm family to a single fold, the model faces a more diverse structural distribution, the prediction task becomes harder, and the metrics deteriorate. This comparison confirms that an algorithm-family-based split is essential for evaluating the true generalisation capability of the model.
Future work could explore representation-learning techniques that encourage structure-invariant features, for example by adding contrastive losses that penalise predictions driven by superficial circuit patterns.
Another aspect worth discussing is the use of simulator-derived execution times as additional features for predicting real-device performance. For the circuits in our dataset with measured execution times on both the simulators and real quantum computers, we measure the Spearman’s rank correlation between each pair of simulators and real quantum computers. Our correlation study (Table 15) shows that, although simulator and hardware run times differ in absolute terms, their relative ordering is highly consistent: Spearman’s rank coefficients range from 0.788 to 0.858 with $p<2.2\times 10^{-16}$. This strong monotonic relationship indicates that simulator timings already provide reliable ranking signals for workload selection and scheduling. Leveraging these signals in hybrid predictive models, perhaps together with simulator back-ends calibrated to capture hardware-specific properties, is therefore a promising direction for future work.
Table 15. Spearman’s rank correlation between the circuits’ execution times on simulators and real quantum computers
Spearman’s Rank Correlation Coefficient Results
FakeWashington vs ibm_osaka $p$-value $2.2\times 10^{-16}$; $\rho=0.825$
FakeWashington vs ibm_kyoto $p$-value $2.2\times 10^{-16}$; $\rho=0.788$
FakeSherbrooke vs ibm_osaka $p$-value $2.2\times 10^{-16}$; $\rho=0.858$
FakeSherbrooke vs ibm_kyoto $p$-value $2.2\times 10^{-16}$; $\rho=0.809$
8. Threats to Validity
8.1. Internal Validity
When executing quantum circuits on quantum computers, the scarcity of quantum computer resources and the lengthy total time required for executing each circuit (including queuing time which usually takes hours for a circuit), prevent us from executing all the quantum circuits on quantum computers. Instead, we selected a subset of the quantum circuits with 95% confidence for execution and data collection. We used active learning to maximize the diversity of the sampled circuits. However, this may result in a deviation between the observed results and the actual results: for example, the performance of our model may increase if more quantum circuits are considered.
8.2. External Validity
@@ 8.3. Construct validity
9. Conclusion
Quantum computing resources are rather limited, thus making judicious planning and utilization of quantum computers is important for both quantum computing service providers and users. In this paper, we first study the characteristics of quantum circuits’ runtime on simulators and real quantum computers. Then, we introduce an innovative graph transformer-based model, utilizing the graph information and global information of quantum circuits to estimate their execution times. Our experiments indicate that the execution times of quantum circuits vary significantly across different circuits, and that our approach can accurately predict their execution times on both simulators and quantum computers. We observe that some characteristics of the quantum circuits, such as the numbers of different types of quantum gates, especially two-qubit gates, play important roles in predicting their execution times, yet we also highlight the intrinsic differences in modeling the execution times of quantum circuits executed on simulators and real computers.
@@ 9. Conclusion
Our study also sheds light on further explorations. First, future research may leverage our approach to automatically optimize quantum circuits for reduced execution time. In addtion, future research may extend our study beyond the IBM Quantum Platform to include other emerging quantum computing technologies, such as Google’s Sycamore or Rigetti, which could provide broader insights for understanding quantum execution times and our model’s adaptability and efficiency across different quantum systems.
Acknowledgment
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) (ALLRP 580933 - 22). We also thank IBM and the Digital Research Alliance of Canada for providing quantum computing and cluster resources for our experiments.
Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) (ALLRP 580933 - 22).